A 50kg object is moving at 18.2 m/s when a 200 N force is applied in the opposite direction of it’s motion. If the object slows down to 12.6 m/s due to the force, how long was the force applied?


1
Expert's answer
2021-06-08T09:34:16-0400

According to the work-energy theorem, the change in the kinetic energy is equal to the work of the force applied to the object:


"\dfrac{mv_2^2}{2} - \dfrac{mv_1^2}{2} = W"

where "m = 50kg" is the mass of the object, "v_1 = 18.2 m\/s, v_2 = 12.6m\/s" are its initial and final speeds respectively, "W = Fs" is the work done by the force "F = 200N" acting along the distance "s". Expressing "s", find:


"\dfrac{mv_2^2}{2} - \dfrac{mv_1^2}{2} = Fs\\
s = \dfrac{m}{2F}(v_2^2 - v_1^2)"

On the other hand, the distance traveled under a constant force (constant acceleration) is given as follows:


"s = v_1t - \dfrac{at^2}{2}"

where "a = \dfrac{F}{m}" is the acceleration of the body (from the second Newton's law) and "t" is the time of the motion. Combining all together, obtain:


"\dfrac{m}{2F}(v_2^2 - v_1^2) = v_1t - \dfrac{Ft^2}{2m}\\
\dfrac{Ft^2}{2m} - v_1t + \dfrac{m}{2F}(v_2^2 - v_1^2) = 0"

Substituting the numbers and solving the quadratic equation for "t", obtain:


"2t^2 - 18.2t - 21.56 = 0\\
t_1 = -1.06s, t_2 = 10.2s"

Taking the positive value (since time can not be negative), obtain the final answer.


Answer. 10.2s.